# A representational framework for learning and encoding structurally enriched trajectories in complex agent environments

### v1

@@ 5 Discussion
In this paper, we have introduced Structurally Enriched Trajectories (SETs), a structured extension of traditional trajectory representations that captures multi-level dependencies between states, interactions, and affordances. SETs are proposed as a fundamental component of a novel computational framework for representation learning, Structurally Enriched Trajectory Learning and Encoding (SETLE), that models hierarchical task structures using a heterogeneous graph memory. By structuring trajectories as SET subgraphs, we extend HeCo’s graph embedding approach to episodic Reinforcement Learning (RL), treating each episode as a distinct subgraph and encoding entire trajectories as central nodes. Our method was able to cluster similar trajectories, constructing a task-focused latent space that can improve cross-task generalisation. Unlike traditional flat trajectory encodings, we provide a way to capture multi-level relationships using network schema and meta-path views, making it particularly effective for complex RL tasks requiring structured trajectory learning. These contributions have broader implications for representation learning, task generalisation, and lifelong AI systems, such as facilitating cross-task adaptability, refining the efficiency of learning from limited experience, and strengthening memory retention for continuous adaptation to new environments.
@@
A key challenge in RL is to ensure that learnt representations transfer across tasks rather than remaining task-specific (D’Eramo et al.,, 2024; Chen et al.,, 2024). By incorporating hierarchical trajectory representations, we enable the recognition of shared patterns and support knowledge transfer across structurally similar settings. Hierarchical structuring is an essential mechanism for generalisation (Sancho et al.,, 2024). In human cognition, for example, learning and adaptation depend on the ability to identify and recombine meaningful substructures across experiences (Lake et al.,, 2017). Many AI models, however, struggle with compositional flexibility, limiting their adaptability to novel tasks and environments. Our framework advances an alternative approach, adopting a hierarchical graph-based representation to model task execution more effectively. This method enables agents to retain and reuse structured knowledge rather than discarding prior experiences, improving adaptation and transfer across tasks.
@@
Although hierarchical structures have been central to reinforcement learning (RL) research for a long time (Botvinick et al.,, 2009), most current approaches depend on predefined skill segmentation. This reliance limits the ability to autonomously discover reusable structures. In contrast, we have modelled the relational dependencies between the trajectory components to ensure that abstract knowledge is connected to the execution-level details. This method aligns with broader trends in structured representation learning (Hafner et al.,, 2019; Roads and Love,, 2024), where explicitly modelling these relationships enhances transferability. By incorporating hierarchical trajectory representations into a structured memory framework, our approach offers a foundation for developing more flexible and scalable reinforcement learning models.
@@
In addition to generalising tasks, this work also advances lifelong learning and adaptive AI. Effective open-ended learning depends on a structured memory system that enables agents to accumulate and refine knowledge over time, rather than starting over with each new task (Finn et al.,, 2017). The hierarchical memory structure in SETLE lays the groundwork for metalearning and continual learning. It allows agents to retain relevant information, adapt to new contexts, and dynamically refine their decision-making strategies. This structured approach also mitigates the issue of catastrophic forgetting, where agents can overwrite previously learnt information.
@@
This work contributes to the broader discussion on structured intelligence in AI (Mohan et al.,, 2024; Eckstein and Collins,, 2020) by providing an effective mechanism to link low-level state-action execution with high-level conceptual reasoning. Achieving human-like adaptability requires us to move beyond mere pattern recognition; it is essential to extract, manipulate, and reuse structured representations across various domains. This exercise in human-level adaptation necessitates continuous interaction among different organisational levels of information and a well-defined hierarchical control structure (Schneider and Logan,, 2006; Hommel et al.,, 2000; Wu et al.,, 2022).
By demonstrating how graph-based trajectory modelling improves task representation, this study lays the groundwork for further exploration of structured reinforcement learning (RL) representations. Hierarchical models have played a crucial role in cognitive science and neuroscience, offering insights into how humans integrate multi-level representations to guide learning and decision-making (Eckstein and Collins,, 2020). While similar approaches in reinforcement learning have been proposed to address long-term dependencies and causal structure learning, challenges still exist in designing frameworks that can generalise beyond task-specific representations (Mohan et al.,, 2024).
While our framework offers several advantages, there are still a number of challenges that need to be addressed. For example, scalability in highly complex environments with non-linear interactions requires further research. Additionally, although hierarchical representations provide a detailed view of trajectories, their efficiency in high-dimensional tasks remains an unresolved issue. Furthermore, this work primarily focusses on representation learning, but understanding its impact on policy optimisation is a key area for future research. Moving forward, exploring how structured trajectory representations can enhance decision-making efficiency will be crucial.

### v2

@@ 6 Discussion
In this paper, we have introduced Structurally Enriched Trajectories (SETs), a structured extension of traditional trajectory representations that captures multi-level dependencies between states, interactions, and affordances. These SETs form the basis of SETLE, a computational framework that extends graph embedding approaches like HeCo to episodic Reinforcement Learning (RL) by structuring trajectories as hierarchical subgraphs. Our experimental results demonstrate that this method successfully constructs a task-focused latent space by clustering similar trajectories. For instance, the encoder achieved a high Silhouette Score of 0.90 for unsuccessful trajectories (with a margin of 1.5) and a low Davies-Bouldin Index of 0.15, indicating that the embeddings formed clear and well-separated clusters based on task outcome. These contributions have broader implications for representation learning, task generalisation, and lifelong AI systems.
The practical advantages of these structured representations were validated when integrated into a full reinforcement learning loop. In complex, physically-grounded tasks, the SETLE-enhanced agent outperformed the baseline; for example, the adapter and penalty strategy achieved a maximum reward over 10.0, while the baseline peaked below 1.0 (Fig. 27). Furthermore, in sparse-reward symbolic environments, the framework fostered more realistic value estimation. The SETLE agent’s Q-values remained stable around 1.0, thereby avoiding the ungrounded value inflation to over 3.0 observed in the baseline agent (Fig. 28).
@@
A key challenge in RL is to ensure that learnt representations transfer across tasks rather than remaining task-specific (D’Eramo et al.,, 2024; Chen et al.,, 2024). By incorporating hierarchical trajectory representations, we enable the recognition of shared patterns and support knowledge transfer. This aligns with findings that hierarchical structuring is an essential mechanism for generalisation (Sancho et al.,, 2024) and reflects principles from human cognition, where learning depends on recombining meaningful substructures from experience (Lake et al.,, 2017). Many AI models, however, struggle with compositional flexibility, limiting their adaptability. The importance of SETLE’s hierarchical structure was empirically validated; our ablation studies demonstrated that flattening the graph representation substantially degraded clustering quality, with the Dunn Index for success clusters dropping from 1.652 to 1.326 (Table 3). This shows that our framework advances a more effective, hierarchical approach that enables agents to retain and reuse structured knowledge.
@@
Although hierarchical structures have been central to RL research for a long time (Botvinick et al.,, 2009), most current approaches depend on predefined skill segmentation, which limits the ability to autonomously discover reusable structures. In contrast, by modelling the relational dependencies between trajectory components, our method aligns with broader trends in structured representation learning (Hafner et al.,, 2019; Roads and Love,, 2024), where explicitly modelling these relationships enhances transferability.
@@
In addition to generalising tasks, this work also advances lifelong learning and adaptive AI. Effective open-ended learning depends on a structured memory system that enables agents to accumulate and refine knowledge over time (Finn et al.,, 2017). The hierarchical memory structure in SETLE lays the groundwork for such capabilities. This adaptability extends across domains; while object-level embeddings failed to align between the CREATE and MiniGrid environments, SETLE’s trajectory-level representations revealed functional similarities, such as a mean cosine similarity of 0.34 between CreateLevelBelt-v0 and MiniGrid-Empty-5x5-v0 (Fig. 35), demonstrating that the framework learns transferable task abstractions. This structured approach also mitigates the issue of catastrophic forgetting.
@@
This work contributes to the broader discussion on structured intelligence in AI (Mohan et al.,, 2024; Eckstein and Collins,, 2020) by providing an effective mechanism to link low-level state-action execution with high-level conceptual reasoning. Achieving human-like adaptability requires moving beyond mere pattern recognition to extract, manipulate, and reuse structured representations across various domains (Schneider and Logan,, 2006; Hommel et al.,, 2000; Wu et al.,, 2022). While similar approaches in RL have been proposed, challenges still exist in designing frameworks that can generalise beyond task-specific representations (Mohan et al.,, 2024).
While our framework offers several advantages, a number of challenges remain. One key limitation is scalability; future work should evaluate SETLE in more complex, real-world environments such as robotics simulations, which involve richer dynamics and higher-dimensional state representations. This, however, poses significant challenges, as the computational complexity of encoding such graphs can be prohibitive, and the efficiency of hierarchical representations in these tasks is an unresolved issue. Additionally, this work primarily focuses on representation learning. A key area for future research is to deepen the understanding of how these structured representations impact policy optimisation. Moving forward, exploring how SETLE can be extended toward dynamic adaptation, enabling agents to continuously refine their representations in response to evolving task demands, will be crucial for pushing reinforcement learning toward lifelong, generalisable intelligence.